# Weak Fano threefolds arising as the blowup of a hyperquadric in $\mathbb{P}^{4}$ along a curve

### v1

@@ 1. Introduction
Replacing $\mathbb{P}^{3}$ or a smooth hypercubic in $\mathbb{P}^{4}$ by a smooth hyperquadric in $\mathbb{P}^{4}$, we obtain similar geometric obstructions on the curve $C$ as in [BL15] and [BL12]. One peculiarity of our case, which does not appear in the other two papers, is that weak Fano threefolds arising from the blowup of a hyperquadric along curves of the same degree and genus can exhibit very different behaviors, depending on the hypersurfaces in which the curves are contained.
Namely, a smooth rational quartic curve $C\subseteq\mathbb{P}^{4}$ can either give rise to a Fano threefold yielding a Sarkisov link of type II if $C$ is not contained in a hyperplane of $\mathbb{P}^{4}$; or otherwise, $C$ gives rise to a weak Fano threefold that is not Fano but yields no Sarkisov link. In a similar manner, the blowup of a smooth hyperquadric $Y\subseteq\mathbb{P}^{4}$ along a smooth irreducible curve $C\subseteq\mathbb{P}^{4}$ of degree 8 and genus 3 either gives rise to a weak Fano threefold not yielding any Sarkisov link if $C$ is contained in a hyperquadric section of $Y$; or otherwise, $C$ gives rise to a weak Fano threefold yielding a Sarkisov link of type II.
@@ 2.1. Arithmetic genus of (possibly singular) curves
If $Z$ is a connected and reduced projective variety (e.g. if $Z$ is integral), recall that $\mathcal{O}_{Z}=\mathbf{k}$ and thus, $\dim_{\mathbf{k}}(H^{0}(Z,\mathcal{O}_{Z}))=1$ (see e.g. [Liu02, Cor. 3.3.21]). MoreoverLet $C\subseteq Y$ be a smooth irreducible curve of genus $g$ and degree $d$ that lies on a smooth hyperquadric $Y\subseteq\mathbb{P}^{4}$ $Z$ is a (possibly non-reduced, singular or reducible) complete intersection in $\mathbb{P}^{n}$ of positive dimension for some $n\geq 1$, then we also have $\dim_{\mathbf{k}}(H^{0}(Z,\mathcal{O}_{Z}))=1$ ([Vak25, 18.6.U]). In particularLet $C\subseteq Y$ be a smooth irreducible curve of genus $g$ and degree $d$ that lies on a smooth hyperquadric $Y\subseteq\mathbb{P}^{4}$ $Z$ is a projective curve that is either connected and reduced or a complete intersection, then the arithmetic genus of $Z$ is just given by $p_{a}(Z)=\dim_{\mathbf{k}}(H^{1}(Z,\mathcal{O}_{Z}))$.
@@ 2.2. Genus of curves on hypersurfaces of a given degree
OtherwiseLet $C\subseteq Y$ be a smooth irreducible curve of genus $g$ and degree $d$ that lies on a smooth hyperquadric $Y\subseteq\mathbb{P}^{4}$ $\ell(K_{C}-D)>0$, then the divisor $D$ is a special effective divisor. Using Clifford’s theorem ([Har77, Thm. IV.5.4]), we obtain $\ell(D)\leq\frac{1}{2}\deg(D)+1=\frac{1}{2}ed+1$, which gives
@@
If $r\in\{0,\>1\}$, then $Y$ is reducible and contains the plane $V_{\mathbb{P}^{4}}(x_{0},\>x_{1})$. If $r=2$, then the plane $V_{\mathbb{P}^{4}}(x_{0}-\omega x_{1},\>x_{2})$ is contained in the singular quadric $Y=V_{\mathbb{P}^{4}}(x_{0}^{2}+x_{1}^{2}+x_{2}^{2})$, where $\omega\in\mathbf{k}$ with $\omega^{2}=-1$. SimilarlyLet $C\subseteq Y$ be a smooth irreducible curve of genus $g$ and degree $d$ that lies on a smooth hyperquadric $Y\subseteq\mathbb{P}^{4}$ $r=3$, then $Y=V_{\mathbb{P}^{4}}(x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2})$ is singular and contains the plane $V_{\mathbb{P}^{4}}(x_{0}-\omega x_{1},\>x_{2}-\omega x_{3})$. The case $r=4$ is the only one for which $Y$ is smooth.
@@ 3. Curves contained on a smooth hypercubic section
3.1. Properties of the anticanonical linear system $|-K_{X}|$
@@
MoreoverLet $C\subseteq Y$ be a smooth irreducible curve of genus $g$ and degree $d$ that lies on a smooth hyperquadric $Y\subseteq\mathbb{P}^{4}$ $X$ is weak Fano, then $|-K_{X}|$ is base-point-free if and only if $|-K_{X}|_{T}|$ is base-point-free (where the restriction $K_{X}|_{T}$ is seen as the restriction of line bundles).
@@ 3.3. Bounding the genus of curves on a smooth hypercubic section
Let $S\subseteq\mathbb{P}^{4}$ be a smooth K3-surface, $D\in\textup{Pic}(S)$ a divisor and $H_{S}$ a general hyperplane section of $S$.
@@
    If $d:=H_{S}\cdot D\geq 0$ and $D^{2}>-4$, then $D\subseteq S$ is an effective (but potentially non-reduced or reducible) curve of degree $d$ and genus $\frac{1}{2}D^{2}+1$.
@@
  • $D=L_{1}+L_{2}+L_{3}$ for three distinct lines $L_{1},L_{2},L_{3}\subseteq S$: Analogously to the second case, we can argue that $L_{i}\cdot L_{j}\leq 1$ for all $i,j\in\{1,2,3\}$ with $i\neq j$. We want to show that $L_{1}\cdot L_{2}+L_{1}\cdot L_{3}+L_{2}\cdot L_{3}\leq 2$.
    Suppose to the contrary that $L_{i}\cdot L_{j}=1$ for all $i,j\in\{1,2,3\}$ with $i\neq j$. Then $L_{1},L_{2}$ and $L_{3}$ intersect in 3 distinct points $p_{1},p_{2},p_{3}\in S$. Consider a plane $H\cap H^{\prime}$ through $p_{1},p_{2}$ and $p_{3}$ that is given as an intersection of two hyperplanes $H,H^{\prime}\subseteq\mathbb{P}^{4}$. Since two points of the set $\{p_{1},p_{2},p_{3}\}\subseteq H\cap H^{\prime}$ are contained in $L_{i}$, we apply Bézout’s theorem to deduce that $L_{i}\subseteq H\cap H^{\prime}$ for all $i\in\{1,2,3\}$. As the smooth hyperquadric $Y$ cannot contain the plane $H\cap H^{\prime}$ (Lemma 2.5(i)), we can again use Bézout’s theorem to argue that $H\cdot H^{\prime}\cdot Y\leq 2$. However, this contradicts the fact that $L_{1}\cup L_{2}\cup L_{3}\subseteq H\cap H^{\prime}\cap Y$.
@@
Moreover, our bounds from 3.13 coincide with the bounds given in [dC97, Thm. 1.4] for all $d\leq 18$, except for the case $d=14$: [dC97, Thm. 1.4] gives $g=15$ as an upper bound. Our bound, however, is $B(14)=16$. This is sharp as we will show in 3.24 (see also 3.25).
@@ 3.4. Curves on a smooth hypercubic section yielding a weak Fano threefold
$g$ $d$ $n$ $C\cdot\Gamma$ $D$ $H\cdot D$ $D^{2}$ $C\cdot D$ $\Gamma\cdot D$
@@
    For instance, 3.24 implies that there is a smooth hyperquadric $Y\subseteq\mathbb{P}^{4}$ admitting a smooth hypercubic section $S$ with the following properties: $S$ contains an irreducible smooth curve $C$ of degree 9 and genus 0 and satisfies $\textup{Pic}(S)=\mathbb{Z}C\oplus\mathbb{Z}H_{S}$, where $H_{S}$ denotes a general hyperplane section of $S$ . $C$ cannot have 4-secant lines nor 7-secant cubics since any such secant-curve would have to be contained in $S$ by Bézout’s theorem. But $S$ does not contain any line or integral conic since $d=9\equiv_{6}3$ (3.23).
@@
$\begin{matrix}\mathbb{P}^{1}&\hookrightarrow&\mathbb{P}^{3}&\hookrightarrow&\mathbb{P}^{4}\\ [s:t]&\mapsto&[s^{3}:s^{2}t:st^{2}:t^{3}]&\mapsto&[s^{3}:s^{2}t:st^{2}:t^{3}]\end{matrix}.$
The line $L:=V_{\mathbb{P}^{4}}(u,v,y)\subseteq\mathbb{P}^{4}$ does not intersect $K$. We can see immediately that $L\cup K$ vanishes on the hypercubic $Z:=V_{\mathbb{P}^{4}}(f)\subseteq\mathbb{P}^{4}$ with $f:=z(u^{2}-vz+xy)+y(x^{2}-vy)$, as well as on the two hyperquadrics $Y:=V_{\mathbb{P}^{4}}(g),\>Q:=V_{\mathbb{P}^{4}}(h)\subseteq\mathbb{P}^{4}$ with $g:=v^{2}-ux-zy$ and h:=vxuyzuh:=vx-uy-zu. Moreover, an explicit calculation shows that the hyperquadric $Y\subseteq\mathbb{P}^{4}$ and the hypercubic section $S:=Y\cap Z$ are smooth.
@@
It follows that $C$ is a smooth irreducible curve of type $(g,d)=(3,8)\in\mathcal{P}$, which admits a 4-secant line $L$ and is contained in a smooth hyperquadric $Y\subseteq\mathbb{P}^{4}$, as well as in a smooth hypercubic section $S$ of $Y$.
@@ 3.5. Existence of curves yielding a weak Fano threefold
Moreover, the following lemma provides conditions for the existence of a rational cubic in $S$. These conditions will be used in 3.24 to show that some of the sextic K3-surfaces of 3.10, which are given by the complete intersection of a quadric and cubic hypersurface of $\mathbb{P}^{4}$, are contained in a smooth hyperquadric.
@@
Recall that all integral conics and cubics in $S$ must be rational since integral cubics in $Y$ must have genus 0 (2.7). By the assumption $H_{1}\cap H_{2}\cap S=3L$, the surface $S$ contains the line $L$. But then $S$ cannot contain an integral conic or cubic by 3.23. This lemma also shows that $L$ is the unique effective divisor in $\textup{Pic}(S)$ of degree 1. For both of these reasons, we must indeed have $R_{1}=3L=R_{2}$.
@@
For $d=14$ and $g=16$, 3.24 proves, in particular, the existence of a smooth hyperquadric $Y\subseteq\mathbb{P}^{4}$, a hypercubic $Z\subseteq\mathbb{P}^{4}$ and a smooth irreducible curve $C$ of degree 14 and genus 16 such that $C$ is contained in the smooth intersection $Y\cap Z$. Since the arithmetic and geometric genus of the smooth curve $C$ coincide, this is a counter example for [dC97, Thm. 1.4], which gives 15 as an upper bound for such a curve.
Now we can prove the main result of this subsection: The existence of a weak Fano threefold arising as the blowup of a smooth hyperquadric in $\mathbb{P}^{4}$ along a smooth irreducible curve $C\subseteq\mathbb{P}^{4}$ whose degree and genus appear in the list $\mathcal{P}$. This result will be a corollary of 3.24 and the following lemma.
@@
Observe that every pair $(g,d)\in\mathcal{P}$ either satisfies $g=\frac{d^{2}}{12}+1$, or $g\leq\frac{d^{2}-1}{12}$ and $(g,d)\neq(4,7)$. So 3.24 gives the existence of a smooth hyperquadric $Y\subseteq\mathbb{P}^{4}$ and a hypercubic $Z\subseteq\mathbb{P}^{4}$ such that the following is satisfied: $S=Y\cap Z$ is a smooth K3-surface that contains a smooth irreducible curve $C$ of degree $d$ and genus $g$ and that satisfies either $\textup{Pic}(S)=\mathbb{Z}H_{S}$, or $\textup{Pic}(S)=\mathbb{Z}H_{S}\oplus\mathbb{Z}C$, where $H_{S}$ denotes a general hyperplane section of $S$.
Then the blowup of $Y$ along $C$ is weak Fano by 3.26.

### v2

@@ 1. Introduction
Replacing $\mathbb{P}^{3}$ or a smooth hypercubic in $\mathbb{P}^{4}$ by a smooth hyperquadric in $\mathbb{P}^{4}$, we obtain similar geometric obstructions on the curve $C$ as in [BL15] and [BL12]. One peculiarity of our case, which does not appear in the other two papers, is that weak Fano threefolds arising from the blowup of a hyperquadric along curves of the same degree $d$ and genus $g$ can exhibit very different behaviors: For some curves of degree $d$ and genus $g$, they give rise to a Sarkisov link, and for other curves of degree $d$ and genus $g$, they do not. This dichotomy appears for the four pairs $(g,d)\in\{(0,4),\>(3,8),\>(6,10),\>(11,12)\}$.
@@ 2.1. Arithmetic genus of (possibly singular) curves
If $Z$ is a connected and reduced projective variety (e.g. if $Z$ is integral), recall that $\mathcal{O}_{Z}=\mathbf{k}$ and thus, $\dim_{\mathbf{k}}(H^{0}(Z,\mathcal{O}_{Z}))=1$ (see e.g. [Liu02, Cor. 3.3.21]). Moreover, $Z$ is a (possibly non-reduced, singular or reducible) complete intersection in $\mathbb{P}^{n}$ of positive dimension for some $n\geq 1$, then we also have $\dim_{\mathbf{k}}(H^{0}(Z,\mathcal{O}_{Z}))=1$ ([Vak25, 18.6.U]). In particular, $Z$ is a projective curve that is either connected and reduced or a complete intersection, then the arithmetic genus of $Z$ is just given by $p_{a}(Z)=\dim_{\mathbf{k}}(H^{1}(Z,\mathcal{O}_{Z}))$.
@@ 2.2. Genus of curves on hypersurfaces of a given degree
Otherwise, $\ell(K_{C}-D)>0$, then the divisor $D$ is a special effective divisor. Using Clifford’s theorem ([Har77, Thm. IV.5.4]), we obtain $\ell(D)\leq\frac{1}{2}\deg(D)+1=\frac{1}{2}ed+1$, which gives
@@
If $r\in\{0,\>1\}$, then $Y$ is reducible and contains the plane $V_{\mathbb{P}^{4}}(x_{0},\>x_{1})$. If $r=2$, then the plane $V_{\mathbb{P}^{4}}(x_{0}-\omega x_{1},\>x_{2})$ is contained in the singular quadric $Y=V_{\mathbb{P}^{4}}(x_{0}^{2}+x_{1}^{2}+x_{2}^{2})$, where $\omega\in\mathbf{k}$ with $\omega^{2}=-1$. Similarly, $r=3$, then $Y=V_{\mathbb{P}^{4}}(x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2})$ is singular and contains the plane $V_{\mathbb{P}^{4}}(x_{0}-\omega x_{1},\>x_{2}-\omega x_{3})$. The case $r=4$ is the only one for which $Y$ is smooth.
@@ 3. Curves contained on a smooth hypercubic section
3.1. Properties of the anticanonical linear system
@@
Moreover, if $X$ is weak Fano, then $|-K_{X}|$ is base-point-free if and only if $|-K_{X}|_{T}|$ is base-point-free (where the restriction $K_{X}|_{T}$ is seen as the restriction of line bundles).
@@ 3.3. Bounding the genus of curves on a smooth hypercubic section
Let $S\subseteq\mathbb{P}^{n}$ be a smooth K3-surface for some $n\in\mathbb{Z}_{>0}$, $D\in\textup{Pic}(S)$ a divisor and $H_{S}$ a general hyperplane section of $S$.
@@
    If $d:=H_{S}\cdot D\geq 0$ and $D^{2}>-4$, then $D\subseteq S\subseteq\mathbb{P}^{n}$ is an effective (but potentially non-reduced or reducible) curve of degree $d$ and genus $\frac{1}{2}D^{2}+1$.
@@
  • $D=L_{1}+L_{2}+L_{3}$ for three distinct lines $L_{1},L_{2},L_{3}\subseteq S$: Analogously to the second case, we can argue that $L_{i}\cdot L_{j}\leq 1$ for all $i,j\in\{1,2,3\}$ with $i\neq j$. We want to show that $L_{1}\cdot L_{2}+L_{1}\cdot L_{3}+L_{2}\cdot L_{3}\leq 2$. Suppose to the contrary that $L_{i}\cdot L_{j}=1$ for all $i,j\in\{1,2,3\}$ with $i\neq j$. Then there are two possibilities:
    The first one is that $L_{1},L_{2}$ and $L_{3}$ all intersect in the same point $p\in S$. If the lines do not lie on a common plane, then the tangent space $T_{p}(S)$ is of dimension at least 3, which contradicts the assumption that $S$ is smooth. Otherwise, if $L_{1},L_{2}$ and $L_{3}$ would lie on a plane $P=H_{1}\cap H_{2}$ for two hyperplanes $H_{1},H_{2}\subseteq\mathbb{P}^{4}$, they must be contained in the intersection $H_{1}\cap H_{2}\cap Q$. However, this intersection must be of degree 2 by Bézout’s Theorem since the smooth hyperquadric $Q$ cannot contain $P$ (2.5). So $H_{1}\cap H_{2}\cap Q$ cannot contain the reducible curve $L_{1}\cup L_{2}\cup L_{3}$ of degree 3.
    The second possibility to have $L_{i}\cdot L_{j}=1$ for all $i,j\in\{1,2,3\}$ with $i\neq j$ is that $L_{1},L_{2}$ and $L_{3}$ intersect in 3 distinct points $p_{1},p_{2},p_{3}\in S$. Consider a plane $H\cap H^{\prime}$ through $p_{1},p_{2}$ and $p_{3}$ that is given as an intersection of two hyperplanes $H,H^{\prime}\subseteq\mathbb{P}^{4}$. Since two points of the set $\{p_{1},p_{2},p_{3}\}\subseteq H\cap H^{\prime}$ are contained in $L_{i}$, we apply Bézout’s theorem to deduce that $L_{i}\subseteq H\cap H^{\prime}$ for all $i\in\{1,2,3\}$. As the smooth hyperquadric $Y$ cannot contain the plane $H\cap H^{\prime}$ (Lemma 2.5(i)), we can again use Bézout’s theorem to argue that $H\cdot H^{\prime}\cdot Y\leq 2$. However, this contradicts the fact that $L_{1}\cup L_{2}\cup L_{3}\subseteq H\cap H^{\prime}\cap Y$.
@@
Moreover, our bounds from 3.13 coincide with the bounds given in [dC97, Thm. 1.4] for all $d\leq 18$, except for the case $d=14$: [dC97, Thm. 1.4] gives $g=15$ as an upper bound. Our bound, however, is $B(14)=16$. This is sharp as we will show in 3.25 (see also 3.26).
@@ 3.4. Curves on a smooth hypercubic section yielding a weak Fano threefold
$g$ $d$ $n$ $C\cdot\Gamma$ $D\in\textup{Pic}(S)$ $H\cdot D$ $D^{2}$ $C\cdot D$ $\Gamma\cdot D$
@@
Remark 3.20.
Let $C\subseteq\mathbb{P}^{3}$ be a smooth irreducible curve of degree $d$ and genus $g$ lying on a smooth quartic surface $S\subseteq\mathbb{P}^{3}$. Then by [BL12], the blowup $X$ of $\mathbb{P}^{3}$ along $C$ is weak Fano if and only if $4d-30\leq g\leq 14$ or $(g,d)=(19,12)$, and $C$ does not admit a 5-secant line, 9-secant conic nor 13-secant twisted cubic. Moreover by [BL12, Prop. 5.8], 13-secant cubics are only possible for $(g,d)\in\{(0,7),\>(2,8),\>(3,8)\}$ and 9-secant conics may only appear for $(g,d)\in\{(6,9),\>(7,9)\}$.
Supposing that there is an irreducible curve $\Gamma\subseteq\mathbb{P}^{3}$ of degree $n\in\{2,3\}$ that is $4n+1$-secant to $C$, we must have $\Gamma\subseteq S$ by Bézout’s theorem. As in the proof of 3.19, we can then try to find an effective divisor $D=aH+bC+c\Gamma\in\mathbb{Z}H\oplus\mathbb{Z}C\oplus\mathbb{Z}\Gamma\subseteq\textup{Pic}(S)$ with $a,b,c\in\mathbb{Z}$ such that
  • $D^{2}=-2$ and $H\cdot D\geq 1$. So $D$ corresponds to an effective rational curve of degree $H\cdot D$ on $S$ by Lemma 3.11(i). (The lemma is applicable since $S\subseteq\mathbb{P}^{3}$ is a smooth K3-surface.)
  • $\deg(D)<d$ and $C\cdot D<0$, or $\deg(D)<n$ and $\Gamma\cdot D<0$, which both contradicts the fact that the intersection number of two effective curves with no common irreducible component is always positive.
  • $H\cdot D=4a+db+nc,\>C\cdot D=da+(2g-2)b+(C\cdot\Gamma)c,\>\Gamma\cdot D=na+(C\cdot\Gamma)b-2c$ and $D^{2}=(H\cdot D)a+(C\cdot D)b+(\Gamma\cdot D)c$.
  • Applying [BL12, Prop. 2.6], we can bound the genus of $C\cup\Gamma$, and thereby the possible intersection numbers $C\cdot\Gamma$, in terms of $g,d$ and $n$ by
[…]
For $(g,d)\in\{(0,7),\>(2,8),\>(3,8),\>(7,9)\}$, we find the contradictory divisor $D\in\textup{Pic}(S)$ as it is given in Table 4. For $(g,d)=(6,9)$, we can only find such a divisor $D$ if $C\cdot\Gamma=10$, but not if $C\cdot\Gamma=9$. In other words, we have shown that the assumptions about 13-secant twisted cubics in [BL12] can be eliminated and that 9-secant conics might only be possible for $(g,d)=(6,9)$.
This is almost the same result as in the paper [D’A17], in which minimal free graded resolutions are used to show that 9-secant conics and 13-secant twisted cubics are never possible and it therefore suffices to assume that no 5-secant lines exist in [BL12].
Table 4. If we suppose that a smooth irreducible curve $C\subseteq\mathbb{P}^{3}$ of degree $d$ and genus $g$ is contained in a smooth quartic surface $S\subseteq\mathbb{P}^{3}$ and that $C$ admits an irreducible curve $\Gamma\subseteq\mathbb{P}^{3}$ of degree $n$ that is at least $4n+1$-secant to $C$, then we can find a contradictory effective divisor $D\in\textup{Pic}(S)$ that has a negative intersection with either $C$ or with $\Gamma$.
$g$ $d$ $n$ $C\cdot\Gamma$ $D\in\textup{Pic}(S)$ $H\cdot D$ $D^{2}$ $C\cdot D$ $\Gamma\cdot D$
0 7 3 13 $H+C+\Gamma 2 -2$ -3 5
0 7 2 $m\in\{9,10,11\}$ $H-\Gamma$ 2 -2 7m7-m 4
2 8 3 14 $3H-C-\Gamma$ 1 -2 8 -3
2 8 3 13 $2H+C+\Gamma$ 3 -2 -1 5
2 8 2 $m\in\{9,10,11\}$ $H-\Gamma$ 2 -2 8m8-m 4
3 8 3 13 $3H-C-\Gamma$ 1 -2 7 -2
3 8 2 $m\in\{9,10\}$ $H-\Gamma$ 2 -2 8m8-m 4
6 9 2 10 $H-\Gamma$ 2 -2 -1 4
7 9 2 9 $3H-C-\Gamma$ 1 -2 6 -1
@@
    For instance, 3.25 implies that there is a smooth hyperquadric $Y\subseteq\mathbb{P}^{4}$ admitting a smooth hypercubic section $S$ with the following properties: $S$ contains an irreducible smooth curve $C$ of degree 9 and genus 0 and satisfies $\textup{Pic}(S)=\mathbb{Z}C\oplus\mathbb{Z}H_{S}$, where $H_{S}$ denotes a general hyperplane section of $S$ . $C$ cannot have 4-secant lines nor 7-secant cubics since any such secant-curve would have to be contained in $S$ by Bézout’s theorem. But $S$ does not contain any line or integral conic since $d=9\equiv_{6}3$ (3.24).
@@
$\begin{matrix}\mathbb{P}^{1}&\hookrightarrow&\mathbb{P}^{3}&\hookrightarrow&\mathbb{P}^{4}\\ [s:t]&\mapsto&[s^{3}:s^{2}t:st^{2}:t^{3}]&\mapsto&[s^{3}:s^{2}t:st^{2}:t^{3}:0]\end{matrix}.$
The line $L:=V_{\mathbb{P}^{4}}(u,v,y)\subseteq\mathbb{P}^{4}$ does not intersect $K$. We can see immediately that the polynomials $f:=z(u^{2}-vz+xy)+y(x^{2}-vy)$, $g:=v^{2}-ux-zy$ and h:=vxuyzuh:=vx-uy-zu vanish on $L\cup K$. So $L\cup K$ is contained in the hypercubic $Z:=V_{\mathbb{P}^{4}}(f)\subseteq\mathbb{P}^{4}$, as well as in the two hyperquadrics $Y:=V_{\mathbb{P}^{4}}(g),\>Q:=V_{\mathbb{P}^{4}}(h)\subseteq\mathbb{P}^{4}$. Moreover, an explicit calculation shows that the hyperquadric $Y\subseteq\mathbb{P}^{4}$ and the hypercubic section $S:=Y\cap Z$ are smooth.
@@
It follows that $C$ is a smooth irreducible curve of type $(g,d)=(3,8)\in\mathcal{P}$, which admits a 4-secant line $L$ and is contained in a smooth hyperquadric $Y\subseteq\mathbb{P}^{4}$, as well as in a smooth hypercubic section $S$ of $Y$. Note that $C$ does not admit a 7-secant conic since $n_{\max}=1$ in 3.19.
@@
Note that $C$ does not admit a 4-secant line since such a line would have to be contained in the intersection $Y\cap Z_{1}\cap Z_{2}$ and then its vanishing ideal would have to appear in the primary decomposition of $I$.
@@ 3.5. Existence of curves yielding a weak Fano threefold
Moreover, the following lemma provides conditions for the existence of a rational cubic in $S$. These conditions will be used in 3.25 to show that some of the sextic K3-surfaces of 3.10, which are given by the complete intersection of a quadric and cubic hypersurface of $\mathbb{P}^{4}$, are contained in a smooth hyperquadric.
@@
Recall that all integral conics and cubics in $S$ must be rational since integral cubics in $Y$ must have genus 0 (2.7). By the assumption $H_{1}\cap H_{2}\cap S=3L$, the surface $S$ contains the line $L$. But then $S$ cannot contain an integral conic or cubic by 3.24. This lemma also shows that $L$ is the unique effective divisor in $\textup{Pic}(S)$ of degree 1. For both of these reasons, we must indeed have $R_{1}=3L=R_{2}$.
@@
For $d=14$ and $g=16$, 3.25 proves, in particular, the existence of a smooth hyperquadric $Y\subseteq\mathbb{P}^{4}$, a hypercubic $Z\subseteq\mathbb{P}^{4}$ and a smooth irreducible curve $C$ of degree 14 and genus 16 such that $C$ is contained in the smooth intersection $Y\cap Z$. Since the arithmetic and geometric genus of the smooth curve $C$ coincide, this is a counter example for [dC97, Thm. 1.4], which gives 15 as an upper bound for such a curve.
Now we can prove the main result of this subsection: The existence of a weak Fano threefold arising as the blowup of a smooth hyperquadric in $\mathbb{P}^{4}$ along a smooth irreducible curve $C\subseteq\mathbb{P}^{4}$ whose degree and genus appear in the list $\mathcal{P}$. This result will be a corollary of 3.25 and the following lemma.
@@
Observe that every pair $(g,d)\in\mathcal{P}$ either satisfies $g=\frac{d^{2}}{12}+1$, or $g\leq\frac{d^{2}-1}{12}$ and $(g,d)\neq(4,7)$. So 3.25 gives the existence of a smooth hyperquadric $Y\subseteq\mathbb{P}^{4}$ and a hypercubic $Z\subseteq\mathbb{P}^{4}$ such that the following is satisfied: $S=Y\cap Z$ is a smooth K3-surface that contains a smooth irreducible curve $C$ of degree $d$ and genus $g$ and that satisfies either $\textup{Pic}(S)=\mathbb{Z}H_{S}$, or $\textup{Pic}(S)=\mathbb{Z}H_{S}\oplus\mathbb{Z}C$, where $H_{S}$ denotes a general hyperplane section of $S$.
Then the blowup of $Y$ along $C$ is weak Fano by 3.27.